max Xi + 3x2 subject to: -X1 – x2 < -3 -21 + x2 = -1 X1 + 2x2 = 2 X1, x2 > 0Problems. Solve the following problems using the simplex method in the dictionary form. Note that problems 2, 3, and 4, requi
Answer #1

The solution too long, so i wrote it down in my word and pasted it here, You may download the image and view it s it will be clearly visible. Thank you!

Solution: Problem is Max Z = x + 3 x2 subject to - i - 22 -3 Here by = -3<0. so multiply this constraint by -1 to make bı > 0

After introducing slack,surplus,artificial variables Max Z = - A1 - A2 subject to x1 + x2 - Si + A = 3 21 - 22 - S2 + A2 = 1

Negative minimum Z - C; is – 2 and its column index is 1. So, the entering variable is 21. Minimum ratio is 1 and its row ind

- Rz (new) = R3 (old) - R2 (new) Rz(old) = 2 1 2 0 R2 (new) = 1 1 -1 0 R3 (new) = R3 (old) - R2(new) 1 0 3 0 0 1 -1 0 1 Itera

Negative minimum 2 - Cis -2 and its column index is 2. So, the entering variable is c.. Minimum ratio is 0.3333 and its row i

- Rz (new) = R2 (old) + R3 (new) R2 (old) = 11 - -1 0 0 R: (new) = 0.3333 0 1 0 0.3333 0.3333 0 Ry(new) = R (old) + R$ (new)

Negative minimum Z; -C; is -0.3333 and its column index is 4. So, the entering variable is S2 Minimum ratio is 1 and its row

- Rz (new) = R2 (old) +0.6667R3 (new) R2 (old) = 1.3333 1 0 0 -0.6667 0.3333 R: (new) = 10 30 0.6667 x R3 (new) = 0.6667 0 2Since all Z - C; > 0 Hence, optimal solution is arrived with value of variables as: 21 = 2, *2 = 0 Max Z=0 But this solution